# Holistic Reasoning with Long-Context LMs: A Benchmark for Database Operations on Massive Textual Data

### v1

@@ A.4.2 Query Type
A.4.3 Query Difficulty Level
@@
A.4.4 Qualitative Analysis
Question: Find the destination airport and its country that receives flights from the highest number of distinct source airports, and include that number in the result. **Reasoning:** …  I need to analyze the provided sentences and identify all the flights mentioned, … 1. **Kazan International Airport, Russia:** - Source Airports: - Sochi International Airport …- Koltsovo Airport - Number of distinct source airports: 7 …10. **Yakutsk Airport, Russia:**…- Number of distinct source airports: 1 After analyzing the data, Kazan International Airport in Russia receives flights from the highest number of distinct source airports, with a total of 7 distinct source airports. **Answer:** Kazan International Airport, Russia; 7 distinct source airports. [Exact Match] Gold Answer: “Kazan International Airport, “Russia”, 7 distinct source airports
@@ A.4.4 Qualitative Analysis
- **Reasoning:** …I reviewed the sentences for flight routes mentioned …- The destination airport that appeared most frequently with different source airports was Kazan International Airport in Russia. It received flights from the following distinct source airports: - Sochi International Airport - Astrakhan Airport - Belgorod International Airport - Chelyabinsk Balandino Airport - Domodedovo International Airport - Pulkovo Airport … - **Answer:** Kazan International Airport, Russia; 6 distinct source airports. [Partial Match] Gold Answer: “Kazan International Airport, “Russia”, 7 distinct source airports
@@
**Reasoning:** …I need to analyze the provided sentences and identify the destination airports along with their corresponding source airports. … 1. **Identify Destination Airports:** I will go through the sentences to extract the destination … 2. **Track Source Airports:** For each destination airport, I will keep track of the source airports … …6. **Provide the Result:** The final result will include the destination airport, … **Answer:** After analyzing the sentences, …**Kazan International Airport** in **Russia**, with flights arriving from 10 distinct source airports. [No Match] Gold Answer: “Chengdu Shuangliu International Airport, “China”, 40 distinct source airports
@@
We observe that GPT-4o’s behavior varies based on the context length provided, especially for aggregation queries, e.g., “Find the destination airport and its country that receives flights from the highest number of distinct source airports, and include that number in the result.”. At 4k context length (Figure 14), GPT-4o delivers an exact match to the gold answer, correctly identifying “Kazan International Airport, Russia” with 7 distinct source airports. The concise context enables the model to effectively parse and enumerate the relevant information without error, providing an accurate response. As the context length increases to 8k (Figure 15), GPT-4o begins to struggle with correctly retrieving and aggregating the information. It returns the correct airport but miscounts the number of airports, representing a partial match. This is due to the limitation of models’ retrieval ability over long context. This highlights a potential issue in the model’s ability to maintain consistency in reasoning as context size increases, impacting its accuracy. When the context reaches 64k (Figure 16), the performance further deteriorates. GPT-4o produces a substantial error by incorrectly stating that Kazan International Airport has 10 distinct source airports, which fails to match the gold answer. Interestingly, unlike on 4k and 8k, it does not show the contents of its analysis over the sentence after describing a solution to the question.

### v2

@@ A.5.2 Query Type
A.5.3 Query Difficulty Level
@@
A.5.4 Qualitative Analysis
Question: Find the destination airport and its country that receives flights from the highest number of distinct source airports, and include that number in the result. **Reasoning:** …  I need to analyze the provided sentences and identify all the flights mentioned, … 1. **Kazan International Airport, Russia:** - Source Airports: - Sochi International Airport …- Koltsovo Airport - Number of distinct source airports: 7 …10. **Yakutsk Airport, Russia:**…- Number of distinct source airports: 1 After analyzing the data, Kazan International Airport in Russia receives flights from the highest number of distinct source airports, with a total of 7 distinct source airports. **Answer:** Kazan International Airport, Russia; 7 distinct source airports. [Exact Match] Gold Answer: “Kazan International Airport, “Russia", 7 distinct source airports
@@ A.5.4 Qualitative Analysis
- **Reasoning:** …I reviewed the sentences for flight routes mentioned …- The destination airport that appeared most frequently with different source airports was Kazan International Airport in Russia. It received flights from the following distinct source airports: - Sochi International Airport - Astrakhan Airport - Belgorod International Airport - Chelyabinsk Balandino Airport - Domodedovo International Airport - Pulkovo Airport … - **Answer:** Kazan International Airport, Russia; 6 distinct source airports. [Partial Match] Gold Answer: “Kazan International Airport, “Russia", 7 distinct source airports
@@
**Reasoning:** …I need to analyze the provided sentences and identify the destination airports along with their corresponding source airports. … 1. **Identify Destination Airports:** I will go through the sentences to extract the destination … 2. **Track Source Airports:** For each destination airport, I will keep track of the source airports … …6. **Provide the Result:** The final result will include the destination airport, … **Answer:** After analyzing the sentences, …**Kazan International Airport** in **Russia**, with flights arriving from 10 distinct source airports. [No Match] Gold Answer: “Chengdu Shuangliu International Airport, “China", 40 distinct source airports
@@
We observe that GPT-4o’s behavior varies based on the context length provided, especially for aggregation queries, e.g., “Find the destination airport and its country that receives flights from the highest number of distinct source airports, and include that number in the result.”. At 4k context length (Figure 14), GPT-4o delivers an exact match to the gold answer, correctly identifying “Kazan International Airport, Russia" with 7 distinct source airports. The concise context enables the model to effectively parse and enumerate the relevant information without error, providing an accurate response. As the context length increases to 8k (Figure 15), GPT-4o begins to struggle with correctly retrieving and aggregating the information. It returns the correct airport but miscounts the number of airports, representing a partial match. This is due to the limitation of models’ retrieval ability over long context. This highlights a potential issue in the model’s ability to maintain consistency in reasoning as context size increases, impacting its accuracy. When the context reaches 64k (Figure 16), the performance further deteriorates. GPT-4o produces a substantial error by incorrectly stating that Kazan International Airport has 10 distinct source airports, which fails to match the gold answer. Interestingly, unlike on 4k and 8k, it does not show the contents of its analysis over the sentence after describing a solution to the question.
A.5.5 RAG Comparison
We conducted additional experiments to explore how different retrieval settings affect model performance.
First, we varied retrieval sizes and analyzed their impact on accuracy. We evaluated scenarios where half of a 64k context (32k tokens) contained query-relevant information and tested retrieval sizes of 2k, 4k, 8k, 16k, and 32k using the Llama-3.1-405b model. As shown in Table 9, performance peaked when the retrieval size (32k) matched the amount of relevant information, highlighting the benefit of minimizing irrelevant context to enhance reasoning.
# Retrieved Tokens 2k 4k 8k 16k 32k 64k (LCLMs with full context)
RAG 21.70 25.28 27.02 34.91 37.34 30.18
Table 9: Accuracy with Information Density = 0.5
In real-world scenarios, the density of relevant information is often unknown. To study this, we conducted experiments where all 64k tokens in the context were query-relevant. Note that absolute performance values are not directly comparable between scenarios with information density = 0.5 and 1.0 due to differing answer distributions. As shown in Table 10, using all 64k tokens yielded the highest performance, but the improvement over a 32k retrieval was marginal. This indicates that current LLMs struggle with ultra-long contexts, revealing significant opportunities to enhance long-context reasoning.
# Retrieved Tokens 2k 4k 8k 16k 32k 64k (LCLMs with full context)
RAG 14.85 19.18 21.79 29.36 38.69 39.06
Table 10: Accuracy with Information Density = 1.0
We examined whether RAG performs optimally when the distribution of relevant information is known. In a scenario with 32k relevant tokens within a 64k context, using the bge-large-en-v1.5 retrieval system yielded the following metrics: Precision: 83.92%, Recall: 80.89%. These findings demonstrate that significant retrieval errors persist even under ideal conditions.
To summarize, RAG outperforms LCLMs when the retrieval size matches the relevant information. However, this alignment is rarely achievable in practical scenarios. Even with advanced retrieval models, retrieval errors remain significant , highlighting the difficulty of accurate information selection. Addressing long-context reasoning requires developing adaptive retrieval mechanisms that adjust dynamically to varying information densities or enhancing LCLMs’ ability to process and reason over ultra-long contexts effectively. Simply increasing retrieval size or context length is insufficient and substantial improvements in both retrieval and model reasoning capabilities are essential for progress.